If the concentration of ethylene glycol (C2H6O2) in your car radiator is 3.55 M what total volume of solution if we use 684g of the ethylene glycol?
1
Expert's answer
2019-04-23T01:59:23-0400

"C_m= \frac{m}{M \cdot V}"

Cm - molarity or molar concentration of the solution, M

m - mass of the solute, g

M - molar mass of the solute, g/mol

V - volume of the solution, L

"V= \frac{m}{M \cdot C_m}"

"M(C_2 H_6 O_2) \approx 2 \cdot 12 + 6 \cdot 1 + 2 \cdot 16 = 62 \; g/mol"

"V= \frac{684}{62 \cdot 3.55} \approx 3.1 \; L"


Answer: 3.1 L.